OpenStudy (anonymous):

10 ft.=to a fraction lowest terms

OpenStudy (tkhunny):

This is not a good question. Can you provide the entire text of the problem statement?

OpenStudy (unklerhaukus):

somethings missing

OpenStudy (calculusfunctions):

10 ft = fraction?? WHAT??!

OpenStudy (unklerhaukus):

\[10[\text{ft}]=120[\text{in}]\]

OpenStudy (calculusfunctions):

Yes but is that her question? I doubt it.

OpenStudy (calculusfunctions):

Why can't people proof read what they've typed, before being so click post happy.

OpenStudy (unklerhaukus):

never mind

OpenStudy (calculusfunctions):

Do you know what she's trying to ask, @UnkleRhaukus ? You're right in that it probably has to do with conversions, but what exactly?

OpenStudy (unklerhaukus):

how could i know?, something is missing from the question

OpenStudy (calculusfunctions):

I thought you were psychic. LOL I'm just saying that it probably has to do with conversions but it's communicated poorly.

OpenStudy (anonymous):

sorry guys this is a measurement and metric problem and i'ts asking to find 10 in.in a ft and in a fraction form so again i'ts 10 in equals how many ft. in an fraction

OpenStudy (unklerhaukus):

\[12[\text{in}]=1[\text{ft}]\]\[1=\frac{1[\text{ft}]}{12[\text{in}]}\] \[\qquad\qquad10[\text{in}]:1[\text{ft}]\]\[10[\text{in}]\times\frac{1[\text{ft}]}{12[\text{in}]}:1[\text{ft}]\]